OpenStudy (anonymous):

The area of the opening of a rectangular window is to be 143 square feet. If the length is to be 2 feet more than the width, what are the dimensions?

OpenStudy (dangerousjesse):

Write the quadratic polynomial on the left hand side in standard form. Expand out terms of the left hand side: \(x^2+2 x = 143\) Take one half of the coefficient of x and square it, then add it to both sides. Add 1 to both sides: \(x^2+2 x+1 = 144\) Factor the left hand side. Write the left hand side as a square: \((x+1)^2 = 144\) Eliminate the exponent on the left hand side. Take the square root of both sides: \(x+1 = 12\) or \(x+1 = -12\) Look at the first equation: Solve for x. Subtract 1 from both sides: \(x = 11\) or \(x+1 = -12\) Look at the second equation: Solve for x. Subtract 1 from both sides...

OpenStudy (dangerousjesse):

Which leaves you with?

OpenStudy (anonymous):

11 by 13?

OpenStudy (anonymous):

@DangerousJesse